The average number of pages in a novel is 326 with a standard deviation of 24

pages. If a sample of 50 novels is randomly chosen, what is the probability that

the average number of pages in these books is between 319 and 331?


1
Expert's answer
2021-06-02T18:01:50-0400

"\mu=326 \\

\sigma=24 \\

n=50 \\

P(319< \bar{X}<331) = P(\bar{X}<331) -P(\bar{X}<319) \\

= P(Z <\frac{331-326}{24/ \sqrt{50}}) -P(Z< \frac{319-326}{24 / \sqrt{50}}) \\

= P(Z< 1.473) -P(Z< -2.062) \\

= 0.9296 -0.0196 \\

= 0.91"